Opened 11 years ago

Closed 10 years ago

Last modified 3 years ago

#231 closed flight processing (fixed)

GB08/13, flight day 077/2009, Lake Vrynwy

Reported by: jatt Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_13-2009_077_Lake_Vrynwy

Data arrived from ARSF via SATA N on 30th March 2009

Scientific details: see /users/rsg/arsf/arsf_data/2008/ARSF_Applications/GB_and_NL/GB08-13 Science Case.pdf

Priority: a4H

PI: Dr Mathias Disney

Sensors:

  • Eagle (8/12/2009)
  • Hawk (8/12/2009)
  • ALS-II Lidar

Quicklook

TBD

Change History (32)

comment:1 Changed 11 years ago by jatt

  • Owner mggr deleted

comment:2 Changed 11 years ago by jatt

  • Milestone set to 2009 Data processing completion

comment:3 Changed 11 years ago by mark1

  • Owner set to mark1

comment:4 Changed 11 years ago by mark1

Applanix has been processed and copied to arsf directory

comment:5 Changed 11 years ago by mark1

Line 4 and 17 in Hawk have Sync errors. SCT offset required for line 17 not sure about 4 yet, need to compare against Eagle data.

comment:6 Changed 11 years ago by mark1

Eagle line 17 has a sync error - may need SCT offset

comment:7 Changed 11 years ago by mark1

We have access to vector information for this flight.

comment:8 Changed 10 years ago by mggr

  • Type changed from task to flight processing

comment:9 Changed 10 years ago by knpa

  • Priority changed from alpha 5 to alpha 4 high

comment:10 Changed 10 years ago by mggr

  • Component changed from ARSF to Processing: general

comment:11 Changed 10 years ago by knpa

  • Owner changed from mark1 to knpa
  • Status changed from new to assigned

comment:12 Changed 10 years ago by knpa

lidar processing begun

comment:13 Changed 10 years ago by knpa

LIDAR base data is incorrectly labeled as 18th March 2009 whereas the contained data is from the 16th and therefore was not matching up against the flight. Obtaining new basestation data from smartnet.

comment:14 Changed 10 years ago by knpa

Processing and QC complete. Delivery procedure able to proceed.

comment:15 Changed 10 years ago by chrfi

delivery check completed.

banding problem noted, waiting on information to correct or document problem.

comment:16 Changed 10 years ago by knpa

Hyperspectral delivery completed and sent out today

comment:17 Changed 10 years ago by knpa

CORRECTION

above note refers to lidar :)

comment:18 Changed 10 years ago by mark1

  • Owner changed from knpa to mark1
  • Status changed from assigned to new

comment:19 Changed 10 years ago by mark1

SCT and GPT offsets for Eagle and Hawk

Eagle Line GPT SCT Hawk Line GPT SCT
1 0 1.0 1 0 -2.5
2 -0.10 0 2 -0.10 0
3 0 0.05 3 0 0
4 0 0 4 0 0.13
5 0 0.05 5 -0.05 0
6 0.02 0 6 0 0
7 0 0 7 0 0
8 -0.07 0 8 -0.05 0
9 -0.03 0 9 0 0
10 -0.05 0 10 -0.10 0
11 -0.05 0 11 -0.05 0
12 -0.05 0 12 0 0
13 -0.05 0 13 -0.03 0
14 -0.05 0 14 -0.06 0
15 -0.05 0 15 -0.05 0
16 -0.05 0 16 0 0
17 -0.10 0 17 0 0.4
18 -0.07 0 18 -0.07 0
19 0 0 19 0 0
20 -0.05 0 20 -0.07 0
21 -0.05 0 21 -0.03 0
22 0 0 22 0 0
23 0 0 23 0 0
24 -0.05 0 24 -0.05 0
25 -0.05 0 25 -0.05 0
26 -0.05 0 26 -0.05 0
27 -0.03 0 27 -0.05 0
28 -0.02 0 28 -0.05 0

comment:20 Changed 10 years ago by mark1

Eagle and hawk per line boresight values

Eagle Line P R H Hawk Line P R H
1 -0.15 -0.65 0.5 1 0.38 0.24 0.45
2 0 -0.70 0 2 -0.18 0.20 0.05
3 -0.15 -0.35 0 3 -0.18 0.20 0
4 -0.04 -0.65 0 4 -0.18 0.24 0
5 -0.04 -0.4 0 5 0.27 0.27 0
6 -0.19 -0.73 0 6 -0.10 0.23 0.30
7 0.02 -0.47 0 7 -0.09 0.23 0
8 0.19 -0.61 -0.2 8 -0.05 0.30 -0.07
9 0.21 -0.55 -0.10 9 0.13 0.16 -0.10
10 -0.08 -0.61 0 10 0.07 0.29 0
11 0.01 -0.54 0.45 11 -0.08 0.14 0.32
12 0.1 -0.67 0.25 12 -0.11 0.20 0.20
13 0.25 -0.54 0 13 0.27 0.15 -0.03
14 -0.002 -0.63 0.10 14 0.17 0.23 0.05
15 0.24 -0.51 0.17 15 0.07 0.17 0.20
16 0.31 -0.57 0 16 0.07 0.29 0
17 0.48 -0.56 0.20 17 -0.20 0.20 0.20
18 0.19 -0.65 0.30 18 0.12 0.22 0.20
19 -0.23 -0.50 0.16 19 -0.32 0.18 0.12
20 0.34 -0.53 0 20 0.11 0.33 -0.02
21 0.015 -0.60 0.07 21 -0.07 0.09 0.08
22 -0.18 -0.64 0.04 22 -0.28 0.22 -0.05
23 -0.20 -0.50 0 23 -0.29 0.18 0.10
24 0.19 -0.65 0 24 0.28 0.24 0.07
25 0.04 -0.55 0.13 25 0.008 0.14 0.04
26 0.08 -0.60 0 26 -0.06 0.28 0.07
27 0 -0.57 0.08 27 0.10 0.13 0.10
28 -0.11 -0.68 0 28 -0.014 0.20 -0.007

comment:21 Changed 10 years ago by emca

Delivery check of hysperspectral data is complete and all data looks good. Data is ready to deliver.
Delivery directory located at:
~mark1/scratch_space/2009_data/GB08_13-2009_077_Lake_Vrynwy/delivery/

comment:22 Changed 10 years ago by mark1

Have moved the lidar delivery directory from within leica directory and into the main ~arsf project directory. i.e. from

/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB08_13-2009_077_Lake_Vrynwy/leica/delivery

to

/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB08_13-2009_077_Lake_Vrynwy/delivery/

comment:23 Changed 10 years ago by mark1

Photographs have been processed and checked.

comment:24 Changed 10 years ago by mark1

  • Description modified (diff)

Eagle, Hawk and Photography have been sent to PI 8th December 2009

comment:25 Changed 10 years ago by mark1

if the PI asks for boresighting info to orthorectify photography, also offer a redelivery with high accuracy ipas tagging instead of applanix sbet

comment:26 Changed 10 years ago by mark1

rsyncing photo delivery to

/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB08_13-2009_077_Lake_Vrynwy

comment:27 Changed 10 years ago by mark1

  • Component changed from Processing: general to Archiving
  • Owner changed from mark1 to benj
  • Status changed from new to assigned

Everything delivered to PI.

Ready for archiving.

comment:28 Changed 10 years ago by knpa

  • Owner changed from benj to knpa

Archived by DTG (12/01/10)
Preparing for NEODC.

comment:29 Changed 10 years ago by knpa

Tarball created and NEODC informed ready for transfer.

comment:30 Changed 10 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Transferred to NEODC.
Closing ticket.

comment:31 Changed 6 years ago by tipo

Reprocessed Hyperspectral.
New SCTs:

Line Eagle Hawk
1 1 -2.48
2 -0.85 -0.12
3 -0.06 -0.12
4 -0.1 -0.12
5 -0.1 -0.12
6 -0.1 -0.12
7 -0.1 -0.12
8 -0.1 -0.12
9 -0.1 -0.12
10 -0.1 -0.12
11 -0.1 -0.12
12 -0.1 -0.12
13 -0.1 -0.12
14 -0.1 -0.12
15 -0.1 -0.12
16 -0.345 0.9
17 -0.1 0.4
18 -0.1 -0.12
19 -0.1 -0.12
20 -0.1 -0.12
21 -0.1 -0.12
22 -0.1 -0.12
23 -0.1 -0.12
24 -0.1 -0.12
25 -0.1 -0.12
26 -0.1 -0.12
27 -0.1 -0.12
28 -0.1 -0.12

comment:32 Changed 3 years ago by lah

Preparing for archive

  • There is a "copy of hawk" raw directory which contains modified files (corresponding file sizes are different). It is unclear which of these should be deleted or what the modifications are so I have left both. (renamed to copy_of_hawk).
  • Leaving kml overview for now (no 2009 directory on server).

Otherwise, ready to upload to NEODC.

Note: See TracTickets for help on using tickets.